Show that the given function is one-to-one and find its inverse. 1 1 f(x) x + 1
Answer #1

Ginen f(n)= att +(24) = f(a) 1 El 24 +1 = 2% te ayti Tthe The = he Be=he E (1) t = (helt » f is Let y = f(x) = 1 at => y = 1

The solution is given in details